# bibliography-2.bib

### v4

@@ Abstract.
An algebraic set $V\subset\mathbb{R}^{n}$ of dimension $d$ is a $\mathbb{Q}$-nonsingular $\mathbb{Q}$-algebraic set if it can be described, both globally and locally, by polynomial equations with rational coefficients. Set $m:=\max\{n,2d+1\}$, if $V$ is compact, or $m:=n+2d+3$, otherwise. We prove that every nonsingular real algebraic set $V\subset\mathbb{R}^{n}$ with nonsingular algebraic subsets $\{V_{i}\}_{i=1}^{\ell}$ in general position is Nash diffeomorphic to a $\mathbb{Q}$-nonsingular $\mathbb{Q}$-algebraic set $V^{\prime}\subset\mathbb{R}^{m}$ with $\mathbb{Q}$-nonsingular $\mathbb{Q}$-algebraic subsets $\{V^{\prime}_{i}\}_{i=1}^{\ell}$ in general position and the Nash diffeomorphism $h:V\to V^{\prime}$ sends each $V_{i}$ to $V_{i}^{\prime}$. As a byproduct we obtain a $\mathbb{Q}$-algebraic relative version of Nash-Tognoli theorem for compact manifolds. A key result in the proof is the description of $\mathbb{Z}/2\mathbb{Z}$-homological cycles of real Grassmannian manifolds by $\mathbb{Q}$-nonsingular $\mathbb{Q}$-algebraic representatives via the Bott-Samelson resolution of real embedded Schubert varieties.
@@ Introduction
At that point it is natural to ask whether the algebraic models of compact smooth manifolds can be further simplified by requiring that the coefficients of the describing polynomial equations belong to a subfield $k$ of ${\mathbb{R}}$ as small as possible. Since the rationals are the smallest subfield of ${\mathbb{R}}$, the final goal is $k={\mathbb{Q}}$.
@@
The answer is affirmative for $k={\mathbb{R}}_{\mathrm{alg}}$, the field of real algebraic numbers, that is, the smallest real closed field. This follows combining Nash’s theorem and the algebraicity result [CS92, Corollary 3.9] for Nash manifolds over any real closed field by Coste and Shiota. For $k={\mathbb{Q}}$ a complete positive answer is given by Ghiloni and the author in [GSa, Theorem 3.9]. Previously, Ballico and Tognoli [BT92, Theorem 0.1] stated a result which is even stronger but, unfortunately, the proof of this theorem is not complete, so the result is not guaranteed to be true. In Remark 4.3 we outline why some arguments of the classical Nash-Tognoli theorem do not directly extend to obtain polynomial equations over ${\mathbb{Q}}$. In particular, we stress that a different notion of regularity over $\mathbb{Q}$ (see Definition 1.4) with respect of the one proposed in [Tog78, Definition 3, p. 30] and used in [BT92] is required to fill the gaps in the proof. However, we recognize to Ballico and Tognoli the original idea of extending the algebraic approximation techniques to obtain nonsingular equations with coefficients over ${\mathbb{Q}}$ for the algebraic models of compact manifolds.
@@
In the algebraic setting, a similar problem is to simplify the description of homeomorphic models of real algebraic sets in terms of the coefficients of the describing equations. Recently, by means of Zariski equisingular techniques, Parusiński and Rond [PR20] proved that actually every real algebraic subset $V\subset{\mathbb{R}}^{n}$ can be deformed by an strongly equisingular semialgebraic and arcwiseanalytic deformation $h_{t}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n}$ to an algebraic subset $V^{\prime}\subset{\mathbb{R}}^{n}$ defined over ${\mathbb{R}}_{\textnormal{alg}}$. Remarkably, no regularity assumptions on the algebraic set $V$ are necessary. In addition, the polynomial equations describing the above model $V^{\prime}$ can be chosen to have coefficients in ${\mathbb{Q}}$ if the extension of ${\mathbb{Q}}$ obtained by adding the coefficients of the polynomial equations defining $V$ is purely transcendental, see [PR20, Remark 13]. However, in general, the fact that ${\mathbb{Q}}$ is not a real closed field is a crucial difficulty. Indeed, the following open problem proposed by Parusiński [Par21] and motivated by Teissier [Tei90] is widely open:
@@
The aim of this paper is to extend latter version over ${\mathbb{Q}}$ of Nash-Tognoli theorem to a relative setting (see Theorem 4.2) providing a general answer to a relative version of [Par21, Open problem 1, p. 199] in the nonsingular case, that is, to positively answer the following question.
@@
We point out that our Main Theorem below, which positively answer above Relative ${\mathbb{Q}}$-algebraicity problem, is very useful in applications. Two directions for applications are providing further classes of algebraic sets that positively answer the general ${\mathbb{Q}}$-algebraicity problem and ${\mathbb{Q}}$-algebraic representatives of ${\mathbb{Z}}/2{\mathbb{Z}}$-homological cycles on manifolds, see [BT80, AK85, BK89, Kuc11]. More in detail, our Main Theorem below is expected to be deeply used to provide a complete positive answer to the ${\mathbb{Q}}$-algebraicity problem in low dimensions. The 1-dimensional case is just a consequence of [GSa, Theorem 5.1 & Corollary 5.2], in dimension $2$ and $3$ the problem is much more challenging. We refer to [BD81, AK92a] for more details on the topological characterization of polyhedra of dimension up to $3$ admitting an algebraic structure and to [CP00] for necessary conditions on local invariants in all dimensions.
Let us outline the structure of the paper.
@@ 1.2. Fundamental examples
It is well-known that ${\mathbb{E}}_{m,n}$ is a connected $\mathscr{C}^{\infty}$-submanifold of ${\mathbb{R}}^{(m+n)^{2}+m+n}$ of dimension $m(n+1)$.
@@
It is well-known that ${\mathbb{E}}^{*}_{m,n}$ is a connected $\mathscr{C}^{\infty}$-submanifold of ${\mathbb{R}}^{(m+n)^{2}}\times{\mathbb{R}}^{m+n}\times{\mathbb{R}}$ of dimension $m(n+1)$.
@@ 3.2. A review on ${\mathbb{Q}}$-nice ${\mathbb{Q}}$-algebraic sets
4. Relative ${\mathbb{Q}}$-algebraicity problem for nonsingular algebraic sets

### v5

@@ Abstract.
We prove a relative version over ${\mathbb{Q}}$ of Nash-Tognoli theorem, that is: Let $M$ be a compact smooth manifold with closed smooth submanifolds $M_{1},\dots,M_{\ell}$ in general position, then there exists a nonsingular real algebraic set $M^{\prime}\subset{\mathbb{R}}^{n}$ with nonsingular algebraic subsets $M_{1}^{\prime},\dots,M_{\ell}^{\prime}$ and a diffeomorphism $h:M\to M^{\prime}$ such that $h(M_{i})=M_{i}^{\prime}$ for all $i=1,\dots,\ell$ such that $M^{\prime},M_{1}^{\prime},\dots,M_{\ell}^{\prime}$ are described, both globally and locally, by polynomial equations with rational coefficients. In addition, if $M,M_{1},\dots,M_{\ell}$ are nonsingular algebraic sets, then we prove the diffeomorphism $h:M\to M^{\prime}$ can be chosen semialgebraic and the result can be extended to the noncompact case. In the proof we describe also the $\mathbb{Z}/2\mathbb{Z}$-homological cycles of real embedded Grassmannian manifolds by nonsingular algebraic representatives over ${\mathbb{Q}}$ via the Bott-Samelson resolution of Schubert varieties.
@@ Introduction
The ${\mathbb{Q}}$-algebraicity problem
At this point, both for applications and for theoretical reasons, a very interesting problem to investigate is the existence of models of smooth compact manifolds, or even models of singular sets, that can be described by polynomial equations with rational coefficients. It is evident how relevant it would be to succeed in this project for algorithmic algebraic geometry: a systematic and efficient description on ${\mathbb{Q}}$ of algebraic sets, both non-singular and singular, would allow exact calculations. In particular, for the purposes of applications, the description over ${\mathbb{Q}}$ of topological models of algebraic sets in low dimensions would be of particular interest in this project. On the other hand, the existence of normal forms over ${\mathbb{Q}}$ (i.e. over ${\mathbb{Z}}$) up to homeomorphism for real algebraic sets allows the reduction modulo $p$ with potential applications of techniques from arithmetic geometry.
Let us first focus on describing models of compact smooth manifolds and real algebraic sets over $k$, where $k$ denotes a subfield of ${\mathbb{R}}$. In particular, the case $k={\mathbb{Q}}$ corresponds to the previous problem and constitutes the hardest task to achieve since ${\mathbb{Q}}$ is the smallest subfield of ${\mathbb{R}}$. Denote by ${\mathbb{R}}_{\mathrm{alg}}$ the field of real algebraic numbers, or in other words, the real closure of ${\mathbb{Q}}$.
@@
For compact smooth manifolds, above problem for $k={\mathbb{R}}_{\mathrm{alg}}$ has an affirmative answer. This follows combining Nash’s theorem and the algebraicity result [CS92, Corollary 3.9] for Nash manifolds over any real closed field by Coste and Shiota. For $k={\mathbb{Q}}$ a positive answer is given by Ghiloni and the author in [GSa, Theorem 3.9]. Previously, Ballico and Tognoli [BT92, Theorem 0.1] stated a result which is even stronger but, unfortunately, the proof of this theorem is not complete, so the result is not guaranteed to be true. In Remark 4.3 we outline why some arguments of the classical Nash-Tognoli theorem do not directly extend to obtain polynomial equations over ${\mathbb{Q}}$. In particular, we stress that several subtleties arise when looking for equations over ${\mathbb{Q}}$ so a different notion of regularity over $\mathbb{Q}$ (see Definition 1.4) with respect of the one proposed in [Tog78, Definition 3, p. 30] and used in [BT92] is required to fill the gaps in the proof. However, we recognize to Ballico and Tognoli the original idea of extending the algebraic approximation techniques to obtain nonsingular equations with coefficients over ${\mathbb{Q}}$ for the algebraic models of compact smooth manifolds.
@@
A similar topological description over $k={\mathbb{R}}_{\mathrm{alg}}$ holds for singular algebraic sets as well. Indeed, by means of Zariski equisingular techniques, Parusiński and Rond [PR20] proved that every real algebraic subset $V\subset{\mathbb{R}}^{n}$ can be deformed by an strongly equisingular semialgebraic and arc-wise analytic deformation $h_{t}:{\mathbb{R}}^{n}\to{\mathbb{R}}^{n}$ to an algebraic subset $V^{\prime}\subset{\mathbb{R}}^{n}$ defined over ${\mathbb{R}}_{\textnormal{alg}}$. Remarkably, no regularity assumptions on the algebraic set $V$ are necessary. In addition, the polynomial equations describing the above model $V^{\prime}$ can be chosen to have coefficients in ${\mathbb{Q}}$ if the extension of ${\mathbb{Q}}$ obtained by adding the coefficients of the polynomial equations defining $V$ is purely transcendental, see [PR20, Remark 13]. However, in general, the fact that ${\mathbb{Q}}$ is not a real closed field is a crucial difficulty. Indeed, the following open problem proposed by Parusiński [Par21] and motivated by Teissier [Tei90] is widely open:
@@
We outline that both the approaches by Coste-Shiota and Parusiński-Rond make use of the model completeness of the theory of real closed fields. That is the reason why their techniques can not be extended to rational numbers, indeed ${\mathbb{Q}}$ is not a real closed field and the theory of ordered fields is not model complete. Hence, the unique approach which seems to work at the moment in investigating the ${\mathbb{Q}}$-algebraicity problem is the one proposed by Ghiloni and the author in [GSa] and which is further developed in the present paper.
@@
In this paper we extend latter version over ${\mathbb{Q}}$ of Nash-Tognoli theorem to a relative setting (see Theorem 4.2) providing a general answer to a relative version of [Par21, Open problem 1, p. 199] in the nonsingular case, that is, to positively answer the following question.
@@
Further developments and open problems
Our Main Theorem below, which positively answers above Relative ${\mathbb{Q}}$-algebraicity problem, is very useful in applications. Let us outline the two main directions for applications.
${\mathbb{Q}}$-algebraicity problem in low dimensions: Our Main Theorem below is expected to be deeply used to provide a complete positive answer to the ${\mathbb{Q}}$-algebraicity problem in low dimensions. The 1-dimensional case is just a consequence of [GSa, Theorem 5.1 & Corollary 5.2], in dimension $2$ and $3$ the problem is much more challenging. We refer to [BD81, AK92a] for more details on the topological characterization of polyhedra of dimension up to $3$ admitting an algebraic structure and to [CP00] for necessary conditions on local invariants in all dimensions.
${\mathbb{Q}}$-algebraic homology of manifolds: It was shown by Thom [Tho54] that not every ${\mathbb{Z}}/2{\mathbb{Z}}$-homology class of compact smooth manifolds can be represented by a closed smooth submanifold, but it can be represented by the fundamental class of a compact smooth manifold by a smooth mapping. As a consequence, Kucharz [Kuc05] showed that any ${\mathbb{Z}}/2{\mathbb{Z}}$-homology class of a compact Nash manifold can be represented by a semialgebraic arc-symmetric subset. On the other hand, there are many homology classes that can not be represented by algebraic subsets (even singular), see [BD84, AK85, BK89, Ben1]. It is interesting then to see whether those classes that can be represented by real algebraic subsets can be realized by ${\mathbb{Q}}$-algebraic subsets.
Structure of the paper & Main Theorem
Let us outline the structure of the paper:
@@ 1.2. Fundamental examples
It is well-known that ${\mathbb{E}}_{m,n}$ is a connected $\mathscr{C}^{\infty}$ submanifold of ${\mathbb{R}}^{(m+n)^{2}+m+n}$ of dimension $m(n+1)$.
@@
It is well-known that ${\mathbb{E}}^{*}_{m,n}$ is a connected $\mathscr{C}^{\infty}$ submanifold of ${\mathbb{R}}^{(m+n)^{2}}\times{\mathbb{R}}^{m+n}\times{\mathbb{R}}$ of dimension $m(n+1)$.
@@ 3.2. A review on ${\mathbb{Q}}$-nice ${\mathbb{Q}}$-algebraic sets
4. Nash-Tognoli theorem over ${\mathbb{Q}}$ the Relative ${\mathbb{Q}}$-algebraicity problem